OpenStudy (anonymous):

Solving equations involving addition and subtraction. I am SO confused, I took a test and got nothing right... Help Please!

OpenStudy (freckles):

5+x=6 ? like this? or like this 5-x=6?

OpenStudy (anonymous):

something like x-16= -36. Sorry I'm kinda stupid :P

OpenStudy (freckles):

you want to isolate the x you have minus 16 on the side with the x to undo the minus you add

OpenStudy (freckles):

so since you have minus 16 you do plug 16 on both sides

OpenStudy (anonymous):

no 1 is stupid you just had a slow moment

OpenStudy (anonymous):

Oh... That makes way more sense then everyone else says.. Thanks :)

OpenStudy (freckles):

x-16=-36 add 16 to both sides x-16+16=-36+16 we did this because we knew -16+16=0 so you would end up with x+0=-36+16 which will allow us to just write x instead of x+0 on that left hand side x=-36+16

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Oh... That makes way more sense then everyone else says.. Thanks :)

OpenStudy (anonymous):

i gave you a medal

OpenStudy (anonymous):

Thanks Xavier!

OpenStudy (anonymous):

your welcome and my name is kasaan

OpenStudy (future_einstein):

no.... you add the 16 if it says minus you do the opposite to each side to isolate x. so the answer t your example is -20

OpenStudy (freckles):

This is what I did above @Future_Einstein

OpenStudy (freckles):

I added 16 on both sides

OpenStudy (future_einstein):

that was good. but xavier did it wrong

OpenStudy (freckles):

oh I didn't see his workings

OpenStudy (freckles):

so I thought you were disagreeing with me at the same time as agreeing :p

OpenStudy (future_einstein):

what? xD